OpenStudy (anonymous):

A coin is tossed and a die with numbers 1-6 is rolled. What is P (tails and 5)

OpenStudy (anonymous):

so get the probability for each outcome, probability for rolling a 5 = 1/6 (for 6 outcomes, and one possibility probability for coining tails = 1/2 (via same reasoning)

OpenStudy (anonymous):

If both events A and B occur on a single performance of an experiment, this is called the intersection or joint probability of A and B, denoted as . P(A [union] B) If two events, A and B are independent then the joint probability is P(A&B) = P(A) * P(B)

OpenStudy (anonymous):

so i multiply the two outcomes?

OpenStudy (anonymous):

so the probablility will be pretty small

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

for example, if two coins are flipped the chance of both being heads is 1/2 * 1/2 = 1/4

OpenStudy (anonymous):

or in your case 1/2 * 1/6 =0.083333 3 recurring

OpenStudy (anonymous):

im pretty sure 1/2 * 1/6 is 1/12?

OpenStudy (anonymous):

yeah it is, and thats the same as 0.083333333333333